KTeV Shift Summary for DAY 9/18/99 (Sat.) -- delayed

Earl C. Swallow (earls@elmhurst.edu)
Sun, 19 Sep 1999 08:24:27 -0500 (CDT)


Shift Summary for DAY shift, 09/18/99 (Sat.)
(transmission delayed by ECS email server failure)

On Shift: Jason Hamm & Earl Swallow/Ashkan + BobT, RickK, RickC, Arun,
Hogan, Elliot, EdB, .....

Hours of beam available at all intensities: 8.4 (out of 8.5)
Hours of beam available at nominal intensity: 8.4
Hours of E799 data (special runs not included): 0.0
Hours of special runs (laser scans, muon runs...):

Accounting of lost time & intensity:

NM2 Access to reconfigure 90-degree monitor.

========================================================================
Runs ended this shift:

14598 BEAM_799 (with 10" Be absorber @ 7E12)

Tapes brought to FCC this shift:

UP(A,B,C)001-003, UPS001 (runs through 14598)

========================================================================
Accesses during shift:

Reconfigure accidental counters, swap 1440 MF2 controller.

========================================================================
Overview of shift:

Continue E799 start-up activities, including: (1) deadtime studies;
(2) work on neutral beam shape (-Y edge soft) and position;
(3) reconfigure accidental counters; (5) onsplit balancing;
(6) obtaining electrons for HCC work.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
14598 07:00 Shift Begins.
Hyperon trigger studies concluded, and a run to
tape with 10" Be absorber and nominal triggers is
in progress.

07:10 End Run 14598 -- 10" Be absorber Run written to
tape. Tapes taken to FCC to provide access to
electrons for calibration.

14600 07:30 Begin studies of deadtime with various triggers
removed to obtain a clearer picture of deadtime
sources (Run 14600 - no tape) -- KTeV Log XXI,
pages 230-231.

09:20 There have been intermittent "laser rms too wide"
warnings, and this time we see problem with laser
illumination corresponding to 1440 MF2. Cycling the
power on this MF (that's Main Frame, not M... F...)
seems to fix the problem. We initially plan to
replace the controller during access, but the fix
seem to be "permanent," so we decide not to mess
with a good thing.

09:30 Ed looks at data volume issues. Streams 5 & 6 are
dominant. Would need to seek te reduce KQ/BAN
volume in Stream 6 and HCC in Stream 5.

09:50 HP4Si is low on toner and all of the cartridges in
control room have bags torn open and appear to be
USED! Encourage cartridge to hang in there by
shaking it -- OK for now, but need to get new ones
Monday.

10:00 RickC begins work on neutral beam.
Moving primary collimator down 2 mm improves neutral
beam shape on low (-Y) side.

12:00 Turn beam off to begin cool-down for RickC access at
NM2 to reconfigure 90-degree monitor.

12:30 Online Split work goes on while beam is off for
access.

12:45 Ashkan covers shift (until ~14:30) so EarlS can go
KTeV Council Meeting

13:15 Access completed. Beam Returns.
Work on neutral beam position continues.

15:30 Shift ends.
RickC reports that he has secondary neutral beam is
aligned as well as possible from online plots --
see KTeV Logbook XXi, pages 234-237.

========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift

Monday, DAY : Get NEW HP4Si toner cartridges from stockroom and put the
old ones (at least 4 in control room) into the FNAL recycling system.

--
===============================================================================
Earl C. Swallow                           email:  earls@elmhurst.edu
Theophil W. Mueller Professor              (or)   earls@fnal.gov
   and Chairperson
Department of Physics                     Phone:  (630)-617-3577
Elmhurst College                          FAX:    (630)-617-3735
190 Prospect Avenue
Elmhurst, IL   60126-3296                 http://www.elmhurst.edu/~earls
===============================================================================
"Books are lighthouses erected in the dark sea of time."
                                --  from Gargoyles  (The Scrolls of Merlin)